# Lower semicontinuity, Stoilow factorization and principal maps

### v2

@@ 3.1. Lower semicontinuity along sequences with asymptotically finite distortion
Let $\mathbf{E}\colon\Omega\times\mathbb{R}^{2\times 2}\to\mathbb{R}\cup\{+\infty\}$ be a Borel measurable functional that is +$\infty$+ ∞ on $\Omega\times\bigr{(}\mathbb{R}^{2\times 2}\setminus\mathbb{R}^{2\times 2}_{+}\bigr{)}$ and such that $\mathbf{E}(z,\cdot)$ is lower semicontinuous on +$\mathbb{R}^{2\times 2}_{+}$ for each $z\in\Omega$. Assume that
@@
[…]
@@
The conclusion of Theorem 3.1 fails in general without the assumption (3.3), but can be reinstated if formulated in terms of biting convergence as in [29]. Under the assumptions of Theorem 3.1 except (3.3) we have that each subsequence of $(u_{j})$ admits a further subsequence $(u_{j_{k}})$ such that
@@
$\limsup_{j\to\infty}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}_{r}(z_{0})}\!|\textup{D}u_{j}|^{2}\,\mathrm{d}m(z)\leq\frac{\lambda\bigl{(}\overline{\mathbb{D}_{r}(z_{0})}\bigr{)}}{\mathscr{L}^{2}\bigl{(}\mathbb{D}_{r}(z_{0})\bigr{)}},$
@@
(3.7) $\limsup_{r\searrow 0}\limsup_{j\to\infty}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}_{r}(z_{0})}\!|\textup{D}u_{j}|^{2}\,\mathrm{d}m(z)\leq\frac{\mathrm{d}\lambda}{\mathrm{d}\mathscr{L}^{2}}(z_{0})<+\infty$
@@
$\lim_{r\searrow 0}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}_{r}(z_{0})}\!\langle\nu_{z},\Phi\rangle\,\mathrm{d}m(z)=\langle\nu_{z_{0}},\Phi\rangle\in\mathbb{R}.$
@@
where $(u_{j})_{z_{0},r}$ is the integral mean of $u_{j}$ over $\mathbb{D}_{r}(z_{0})$. Hereby $v_{j,r}\in\mathrm{W}^{1,2}(\mathbb{D})$ and $(v_{j,r})_{0,1}=0$. Before the next display we introduce some convenient notation for functions $\theta\colon\mathbb{C}\to\mathbb{R}$: the reflected function, $\tilde{\theta}$, is defined as $\tilde{\theta}(z)\equiv\theta(-z)$, and the $\mathrm{L}^{1}$-dilated function with factor $r>0$, $\theta_{r}$, is defined as $\theta_{r}(z)\equiv\theta\bigl{(}z/r\bigr{)}/r^{2}$. In these terms we note that for each $z_{0}\in\Omega\setminus\mathcal{N}$ and $\eta\otimes\Phi\in\mathcal{F}$ the above Lebesgue point property and a standard result about convolution yield (here $\eta$ is extended to $\mathbb{C}\setminus\mathbb{D}$ by 00):
@@
$\liminf_{j\to\infty}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}}\!\mathbf{1}_{\mathcal{E}}(\textup{D}v_{j})\,\mathrm{d}m(z)\geq\liminf_{j\to\infty}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}}\!\Phi_{k}(\textup{D}v_{j})\,\mathrm{d}m(z)=\langle\nu_{z_{0}},\Phi_{k}\rangle$
@@
$\liminf_{j\to\infty}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}}\!\mathbf{1}_{\mathcal{E}}(\textup{D}v_{j})\,\mathrm{d}m(z)\geq\nu_{z_{0}}\bigl{(}\mathcal{E}\bigr{)}.$
To conclude the proof we estimate the right-hand side using (3.11) for $j\geq j_{a}$:
[…]
@@
and the latter tends by (3.10) to 00 as $j\to\infty$. Consequently, $\nu_{z_{0}}\bigl{(}\mathcal{E}\bigr{)}=0$. Recall that $\mathcal{E}=\mathcal{E}_{a}$ and that $a>1$ is arbitrary, so by monotone convergence we deduce that $\nu_{z_{0}}\bigl{(}\bigcup_{a>1}\mathcal{E}_{a}\bigr{)}=0$. But then $\nu_{z_{0}}$ is carried by $Q_{2}(K(z_{0}))$, and since $Q_{2}(K(z_{0}))$ is a closed set it must contain the support of $\nu_{z_{0}}$, as required. The above holds for all $z_{0}\in G\setminus\mathcal{N}$ that are density points of $G$, where we recall that $G$ was a measurable subset of $\mathbb{A}$ such that $\mathscr{L}^{2}(\mathbb{A}\setminus G)<\varepsilon$. Here $\varepsilon>0$ was arbitrary and it is then routine to conclude the proof. ∎
@@
and that the sequence $(\textup{D}u_{j})$ generates the Young measure $(\nu_{z})_{z\in\Omega}$. Here, for each $z\in\Omega$, the partial functional $\mathbf{E}(z,\cdot)$ is lower semicontinuous on +$\mathbb{R}^{2\times 2}_{+}$, and so it coincides with its lower semicontinuous envelope
@@
Moreover, $\mathbf{E}(z,\cdot)$ is principal quasiconvex and real-valued on +$\mathbb{R}^{2\times 2}_{+}$, so must be rank-one convex and thus locally Lipschitz on +$\mathbb{R}^{2\times 2}_{+}\setminus\{0\}$. It need not be continuous at 00 relative to +$\mathbb{R}^{2\times 2}_{+}$, but is lower semicontinuous there by assumption.
@@
Next, for $\mathscr{L}^{2}$ a.e. $z\in\mathbb{A}$, from [3] we find a sequence of principal $K(z)$-quasiconformal maps $f_{j}\colon\mathbb{C}\to\mathbb{C}$ such that $f_{j}\rightharpoonup\langle\nu_{z},\mathrm{Id}\rangle$ in $\mathrm{W}^{1,p}(\mathbb{D})$ for each $p<2K(z)/(K(z)-1)$ and $(\textup{D}f_{j})$ generates $\nu_{z}$. Recall that according to the 0-1 Law for homogeneous gradient Young measures from [3] we have that either $\nu_{z}(\{0\})=0$ or $\nu_{z}=\delta_{0}$. For measures $\nu_{z}$ of the former type, $\mathbf{E}(z,\cdot)$ is therefore continuous $\nu_{z}$ almost everywhere and so as the quadratic growth (3.1) ensures that the sequence $\bigl{(}\mathbf{E}(\textup{D}f_{j})\bigr{)}$ is equiintegrable on $\mathbb{D}$, we arrive at
@@
For the other measures $\nu_{z}=\delta_{0}$ we simply use that $\mathbf{E}(z,\cdot)$ is lower semicontinuous at 00, that the quadratic growth (3.1) still ensures that the sequence $\bigl{(}\mathbf{E}(\textup{D}f_{j})\bigr{)}$ is equiintegrable on $\mathbb{D}$ (in particular the negative parts) and then the general lower semicontinuity result for Young measures [10]. This concludes the proof. ∎
3.2. Lower semicontinuity with integrable distortion
In this subsection we give the proof of a more general version of Theorem 1.6, which follows the proof in [4, Section 12] concerning the lower semicontinuity of the non polyconvex functional $W$. This functional was introduced in [6] and was also studied in the recent works [46, 47]. As the result below shows, the crucial feature of $W$ is its principal quasiconvexity.
Theorem 3.4.
Let $\mathbf{E}\colon\Omega\times\mathbb{R}^{2\times 2}\to\mathbb{R}\cup\{+\infty\}$ be a Borel measurable functional that is +$\infty$+ ∞ on $\Omega\times\bigr{(}\mathbb{R}^{2\times 2}\setminus\mathbb{R}^{2\times 2}_{+}\bigr{)}$ and such that $\mathbf{E}(z,\cdot)$ is lower semicontinuous on +$\mathbb{R}^{2\times 2}_{+}$ for each $z\in\Omega$ and principal quasiconvex. Assume that there exists constants $C_{1},C_{2}$ such that
(3.12) ${\bf E}^{-}(z,A)\leqslant C_{1}(\max\{|A|^{2-\epsilon},-\log(\det A),K_{A}\}+C_{2}$
Let $g$ be a globally defined homeomorphism in $\mathbb{C}$ with $K_{g}\in L^{1}$ and consider a sequence of maps $u_{j}\in g+\mathrm{W}^{1,2}_{0}(\Omega)$ with $K_{u_{j}}\in L^{q}$ for $q>1$. Then, if $u_{j}\rightharpoonup u$ in $\mathrm{W}^{1,2}_{\operatorname{loc}}(\Omega)$ we have
The proof follows the rough strategy used in the previous subsection. As a first step towards this theorem, we first show that the growth conditions assumed are enough for equiintegrability: this relies on [33, 32] as presented in [4, Proposition 12.8]. By Young measure theory the proof then reduces to obtaining Jensen’s inequality for a homogeneous gradient Young measure, to which Theorem 1.4 then applies. The conclusion then follows from principal quasiconvexity.
Proposition 3.5.
Let ${\bf E}$ satisfy (1.11) and let $(u_{j})\subset\mathrm{W}^{1,2}_{g}(\Omega)$ be a sequence bounded in $\mathrm{W}^{1,2}(\Omega)$, and $\sup_{j}\|K_{u_{j}}\|_{\mathrm{L}^{q}(\Omega)}<\infty$. Then ${\bf E}(\textup{D}u_{j}))$ is equiintegrable in $\Omega$.
Proof.
Recall (1.11) and that by our assumptions both $|\textup{D}u_{j}|^{2-\epsilon}$ and $K_{u_{j}}$ are equiintegrable. It remains to show that $\log J_{u_{j}}$ is equiintegrable but this follows from the pointwise estimate:
$|\log(J_{u_{j}})|\leq\log\left(e+\frac{1}{J_{u_{j}}}\right)+(J_{u_{j}})^{\frac{1}{2}}.$
Namely the first term is directly taken care by [4, Proposition 12.8], and the second term from the boundary values. ∎
Proof of theorem 3.4.
First note that, by the boundary values, the sequence $u_{j}$ consists of homeomorphisms. By passing to subsequences we may assume that $u_{j}\rightharpoonup u$ in $\mathrm{W}^{1,2}(\Omega)$ and, by the fundamental theorem of Young Measures [4, Theorem 2.8], that $u_{j}$ generates the $\mathrm{W}^{1,2}$-gradient Young measure $(\nu_{z})_{z\in\Omega}$. Since $({\bf E}(z,\textup{D}u_{j}))$ is equi-integrable by Proposition 3.5, we have
(3.13) $\lim_{j\to\infty}\int_{\Omega}{\bf E}(z,\textup{D}u_{j})\,\mathrm{d}m(z)=\int_{\Omega}\int_{\mathbb{R}^{2\times 2}}{\bf E}(z,A)\,\mathrm{d}\nu_{z}(A)\,\mathrm{d}m(z).$
For a.e. $z\in\Omega$ we have
(3.14) $\det\overline{\nu}_{z}>0\quad\textup{and}\quad\nu_{z}\in\mathscr{M}^{2}_{\textup{qc}},$
where the first claim follows by considering the functional $\min\{0,\det\}$ and applying it to the Young measure, and the latter claim due to Proposition 3.3.
@@ 3.2. Lower semicontinuity with integrable distortion
Let us fix a point $z\in\Omega$ for which (3.14) holds. The measure $\nu_{z}$ is generated by taking a diagonal subsequence $(\psi_{j})$ of $\psi_{j,\lambda}(w)=\lambda^{-1}u_{j}(z_{0}+\lambda w)$, where $j\to\infty$ and $\lambda\to 0$, cf.  [4, Theorem 2.8]. In particular, $\psi_{j}\colon\mathbb{D}\to\mathbb{C}$ is a sequence of homeomorphisms such that $\sup_{j}\|K_{\psi_{j}}\|_{\mathrm{L}^{q}(\mathbb{D})}<\infty$. Then we are precisely in position to apply to this sequence Theorem 1.4 to find sequences $(h_{j}),(f_{j})$ of respectively conformal and principal maps. Let us call $A_{z}$ the center of mass of $\nu_{z}$, and notice that $f_{j}\to A_{z}$ locally uniformly in $\mathbb{C}$, $\textup{D}f_{j}$ converges weakly to $A_{z}$, $A_{f_{j}}=\int_{\textup{D}}\textup{D}f_{j}\,\mathrm{d}m(z)$ converges to $A_{z}$, and $K_{f_{j}}=K_{\psi_{j}}$ a.e. in $\mathbb{D}$. Therefore as in Proposition 3.5 it follows that ${\bf{E}}(z,\textup{D}f_{j}(w))$ equi-integrable over $\mathbb{D}$. Notice that $f_{j}$ is globally defined and $K_{f_{j}}\equiv 1$ off the unit disc. Hence the proof of Jensen inequality finishes as follows:
$\int_{\mathbb{R}^{2\times 2}}{\bf E}(z,A)\,\mathrm{d}\nu_{z}(A)$ $=\lim_{j\to\infty}\fint_{\mathbb{D}(0,R)}{\bf E}(z,\textup{D}f_{j}(w))\,\mathrm{d}m(w)$
$\to\lim_{j\to\infty}\fint_{\mathbb{D}}{\bf E}(z,,\textup{D}f_{j}(w))\,\mathrm{d}m(w)\geqslant\lim_{j\to\infty}{\bf E}(A_{f_{j}})=E(A_{z})$

### v3

@@ 3.1. Lower semicontinuity along sequences with asymptotically finite distortion
Let $\mathbf{E}\colon\Omega\times\mathbb{R}^{2\times 2}\to\mathbb{R}\cup\{+\infty\}$ be a normal integrand. Assume that
@@
(3.2) $K\equiv\limsup_{j\to\infty}\mathbf{K}(\textup{D}u_{j})<+\infty\quad\mathscr{L}^{2}\mbox{ a.e.~in }\mathbb{A}.$
@@
The conclusion of Theorem 3.1 fails in general without the assumption (3.3), but can be reinstated if formulated in terms of biting convergence as in [51] and [41]. Under the assumptions of Theorem 3.1 except (3.3) we have that each subsequence of $(u_{j})$ admits a further subsequence $(u_{j_{k}})$ such that
@@
$\limsup_{j\to\infty}\fint_{\mathbb{D}_{r}(z_{0})}\!|\textup{D}u_{j}|^{2}\,\mathrm{d}m(z)\leq\frac{\lambda\bigl{(}\overline{\mathbb{D}_{r}(z_{0})}\bigr{)}}{\mathscr{L}^{2}\bigl{(}\mathbb{D}_{r}(z_{0})\bigr{)}},$
@@
(3.7) $\limsup_{r\searrow 0}\limsup_{j\to\infty}\fint_{\mathbb{D}_{r}(z_{0})}\!|\textup{D}u_{j}|^{2}\,\mathrm{d}m(z)\leq\frac{\mathrm{d}\lambda}{\mathrm{d}\mathscr{L}^{2}}(z_{0})<+\infty$
@@
$\lim_{r\searrow 0}\fint_{\mathbb{D}_{r}(z_{0})}\!\langle\nu_{z},\Phi\rangle\,\mathrm{d}m(z)=\langle\nu_{z_{0}},\Phi\rangle\in\mathbb{R}.$
@@
where $(u_{j})_{z_{0},r}$ is the integral mean of $u_{j}$ over $\mathbb{D}_{r}(z_{0})$. Hereby $v_{j,r}\in\mathrm{W}^{1,2}(\mathbb{D})$ and $(v_{j,r})_{0,1}=0$. Before the next display we introduce some convenient notation for functions $\theta\colon\mathbb{C}\to\mathbb{R}:$ the reflected function, $\tilde{\theta}$, is defined as $\tilde{\theta}(z)\equiv\theta(-z)$, and the $\mathrm{L}^{1}$-dilated function with factor $r>0$, $\theta_{r}$, is defined as $\theta_{r}(z)\equiv\theta\bigl{(}z/r\bigr{)}/r^{2}$. In these terms we note that for each $z_{0}\in\Omega\setminus\mathcal{N}$ and $\eta\otimes\Phi\in\mathcal{F}$ the above Lebesgue point property and a standard result about convolution yield (here $\eta$ is extended to $\mathbb{C}\setminus\mathbb{D}$ by 0):
@@
$\liminf_{j\to\infty}\fint_{\mathbb{D}}\!\mathbf{1}_{\mathcal{E}}(\textup{D}v_{j})\,\mathrm{d}m(z)\geq\liminf_{j\to\infty}\fint_{\mathbb{D}}\!\Phi_{k}(\textup{D}v_{j})\,\mathrm{d}m(z)=\langle\nu_{z_{0}},\Phi_{k}\rangle$
@@
$\liminf_{j\to\infty}\fint_{\mathbb{D}}\!\mathbf{1}_{\mathcal{E}}(\textup{D}v_{j})\,\mathrm{d}m(z)\geq\nu_{z_{0}}\bigl{(}\mathcal{E}\bigr{)}.$
To conclude the proof we estimate the left-hand side using (3.11) for $j\geq j_{a}$:
$\fint_{\mathbb{D}}\!\mathbf{1}_{\mathcal{E}}(\textup{D}v_{j})\,\mathrm{d}m(z)$ $=$ $\frac{1}{\pi}\left(\int_{\mathbb{D}\cap(G_{j}-z_{0})/r_{j}}+\int_{\mathbb{D}\setminus(G_{j}-z_{0})/r_{j})}\right)\mathbf{1}_{\mathcal{E}}(\textup{D}v_{j})\,\mathrm{d}m(z)$
@@
and the latter tends by (3.10) to 0 as $j\to\infty$. Consequently, $\nu_{z_{0}}\bigl{(}\mathcal{E}\bigr{)}=0$. Recall that $\mathcal{E}=\mathcal{E}_{a}$ and that $a>1$ is arbitrary, so by monotone convergence we deduce that $\nu_{z_{0}}\bigl{(}\bigcup_{a>1}\mathcal{E}_{a}\bigr{)}=0$. But then $\nu_{z_{0}}$ is carried by $Q_{2}(K(z_{0}))$, and since $Q_{2}(K(z_{0}))$ is a closed set it must contain the support of $\nu_{z_{0}}$, as required. The above holds for all $z_{0}\in G\setminus\mathcal{N}$ that are density points of $G$, where we recall that $G$ was a measurable subset of $\mathbb{A}$ such that $\mathscr{L}^{2}(\mathbb{A}\setminus G)<\varepsilon$. Here $\varepsilon>0$ was arbitrary and it is then routine to conclude the proof. ∎
@@
and that the sequence $(\textup{D}u_{j})$ generates the Young measure $(\nu_{z})_{z\in\Omega}$. Here, for each $z\in\Omega$, the partial functional $\mathbf{E}(z,\cdot)$ is lower semicontinuous on $\mathbb{R}^{2\times 2}_{+}$, and so it coincides with its lower semicontinuous envelope
@@
Moreover, $\mathbf{E}(z,\cdot)$ is principal quasiconvex and real-valued on $\mathbb{R}^{2\times 2}_{+}$, so must be rank-one convex and thus locally Lipschitz on $\mathbb{R}^{2\times 2}_{+}\setminus\{0\}$. It need not be continuous at 0 relative to $\mathbb{R}^{2\times 2}_{+}$, but is lower semicontinuous there by assumption.
@@
Next, for $\mathscr{L}^{2}$ a.e. $z\in\mathbb{A}$, from [4] we find a sequence of principal $K(z)$-quasiconformal maps $f_{j}\colon\mathbb{C}\to\mathbb{C}$ such that $f_{j}\rightharpoonup\langle\nu_{z},\mathrm{Id}\rangle$ in $\mathrm{W}^{1,p}(\mathbb{D})$ for each p<2K(z)/(K(z)1)p<2K(z)/(K(z)-1) and $(\textup{D}f_{j})$ generates $\nu_{z}$. Recall that according to the 0-1 Law for homogeneous gradient Young measures from [4] we have that either $\nu_{z}(\{0\})=0$ or $\nu_{z}=\delta_{0}$. For measures $\nu_{z}$ of the former type, $\mathbf{E}(z,\cdot)$ is therefore continuous $\nu_{z}$ almost everywhere and so as the quadratic growth (3.1) ensures that the sequence $\bigl{(}\mathbf{E}(\textup{D}f_{j})\bigr{)}$ is equiintegrable on $\mathbb{D}$, we arrive at
$\fint_{\mathbb{D}}\!\mathbf{E}(z,\textup{D}f_{j}(w))\,\mathrm{d}m(w)\to\langle\nu_{z},\mathbf{E}(z,\cdot)\rangle.$
For the other measures $\nu_{z}=\delta_{0}$ we simply use that $\mathbf{E}(z,\cdot)$ is lower semicontinuous at 0, that the quadratic growth (3.1) still ensures that the sequence $\bigl{(}\mathbf{E}(\textup{D}f_{j})\bigr{)}$ is equiintegrable on $\mathbb{D}$ (in particular the negative parts) and then the general lower semicontinuity result for Young measures [14]. This concludes the proof. ∎
3.2. Lower semicontinuity with integrable distortion
In this subsection we give the proof of a more general version of Theorem 1.6, which follows the proof in [5, Section 12] concerning the lower semicontinuity of the non polyconvex functional W whose definition was recalled at (2.5). This functional is the Shield transform of a functional introduced in [52], [10] and it was also studied in the recent works [85, 86]. As the result below shows, the crucial feature of W is its principal quasiconvexity.
Theorem 3.4.
Let $\mathbf{E}\colon\Omega\times\mathbb{R}^{2\times 2}\to\mathbb{R}\cup\{+\infty\}$ be a Borel measurable functional such that $\mathbf{E}(z,\cdot)$ is $\mathrm{W}^{1,1}$ principal quasiconvex for each $z\in\Omega$. Assume that there exist constants $C>0$ and $p\in[1,2)$ such that
(3.12) $\bigl{|}{\bf E}(z,A)\bigr{|}\leq C\max\bigl{\{}|A|^{p},-\log(\det A),K_{A}\bigr{\}}+C$
holds for all $(z,A)\in\Omega\times\mathbb{R}^{2\times 2}_{+}$, while ${\bf E}(z,A)=+\infty$ for $(z,A)\in\Omega\times\bigl{(}\mathbb{R}^{2\times 2}\setminus\bigl{(}\mathbb{R}^{2\times 2}_{+}\cup\{0\}\bigr{)}\bigr{)}$. Let $q>\frac{p}{2-p}$ be an exponent and $g\in\mathrm{W}^{1,2}_{\operatorname{loc}}(\mathbb{C})$ a homeomorphism with $K_{g}\in\mathrm{L}^{q}_{\operatorname{loc}}(\mathbb{C})$. If $(u_{j})$ is a sequence of homeomorphisms, $u_{j}\colon\Omega\to u_{j}(\Omega)$, $u_{j}\in g+\mathrm{W}^{1,2}_{0}(\Omega)$ with $\sup_{j}\|K_{u_{j}}\|_{\mathrm{L}^{q}(\Omega)}<+\infty$ and $u_{j}\rightharpoonup u$ in $\mathrm{W}^{1,2}_{\operatorname{loc}}(\Omega)$, then $u\in g+\mathrm{W}^{1,2}_{0}(\Omega)$ is a homeomorphism with $K_{u}\in\mathrm{L}^{q}(\Omega)$ and
Apart from the use of principal quasiconvexity the overall proof strategy is standard. The implementation is however not and it might have other applications and so be of wider interest. First we show that the assumed growth conditions suffice to ensure equiintegrability of the energy densities. This relies on [58, 57] as presented in [5, Proposition 12.8]. Next, by Young measure theory the proof then reduces to obtaining Jensen’s inequality for a homogeneous gradient Young measure, which is the content of Proposition 1.4.
Proposition 3.5.
Let $\bf E$} satisfy (3.12) and let $(v_{j})$ be a bounded sequence in $\mathrm{W}^{1,2}(\Omega)$ with $\sup_{j}\|K_{u_{j}}\|_{\mathrm{L}^{q}(\Omega)}<\infty$ for some $q>\frac{p}{2-p}$. Then the sequence $\bigl{(}{\bf E}(\cdot,\textup{D}v_{j})\bigr{)}$ is equiintegrable in $\Omega$.
Proof.
The assumptions imply that both $|\textup{D}v_{j}|^{p}$ and $K_{v_{j}}$ are equiintegrable. In order to see that $\log J_{v_{j}}$ is equiintegrable too we invoke the pointwise estimate:
(3.13) $|\log(J_{v_{j}})|\leq\log\left(e+\frac{1}{J_{v_{j}}}\right)+(J_{v_{j}})^{\frac{1}{2}}.$
Here the first term on the right-hand side is equiintegrable by virtue of [5, Proposition 12.8], while equiintegrability of the second follows from the assumed $\mathrm{W}^{1,2}$ bound. ∎
Proof of Theorem 3.4.
It is well-known that the limit $u\in g+\mathrm{W}^{1,2}_{0}(\Omega)$ is a homeomorphism with distortion $K_{u}\in\mathrm{L}^{q}(\Omega)$, see [8, Ch. 21]. In view of Proposition 3.5 the sequence $\bigl{(}{\bf E}(\cdot,\textup{D}u_{j})\bigr{)}$ is equiintegrable on $\Omega$, and for a subsequence that we for convenience do not relabel we have
$\int_{\Omega}\!\mathbf{E}(\cdot,\textup{D}u_{j})\,\mathrm{d}m(z)\to\ell\in\mathbb{R}.$
Extracting a further subsequence, again not relabelled, we may assume that $(\textup{D}u_{j})$ generates the $\mathrm{W}^{1,2}$-gradient Young measure $(\nu_{z})_{z\in\Omega}$. First, note that the $\mathrm{L}^{q}$ bound entails
$\int_{\Omega}\!\langle\nu_{z},{\bf K}\rangle\,\mathrm{d}m(z)=\lim_{j\to\infty}\int_{\Omega}\!K_{u_{j}}\,\mathrm{d}m(z)<+\infty,$
so that, in particular, $\langle\nu_{z},{\bf K}\rangle<+\infty$ holds for a.e. $z\in\Omega$. If therefore ${\bf E}_{\ast}(z,\cdot)$ denotes the lower semicontinuous envelope of ${\bf E}(z,\cdot)$, then ${\bf E}_{\ast}(z,\cdot)={\bf E}(z,\cdot)$ $\nu_{z}$-a.e. for $\mathscr{L}^{2}$ a.e. $z\in\Omega$. Hence we get by general Young measure theory [14],
(3.14) $\lim_{j\to\infty}\int_{\Omega}{\bf E}(z,\textup{D}u_{j})\,\mathrm{d}m(z)=\ell\geq\int_{\Omega}\int_{\mathbb{R}^{2\times 2}}{\bf E}(z,A)\,\mathrm{d}\nu_{z}(A)\,\mathrm{d}m(z)>-\infty.$
For a.e. $z\in\Omega$ it is routine to check that we have
@@ 3.2. Lower semicontinuity with integrable distortion
Let us fix a point $z\in\Omega$ for which (3.15) holds. The measure $\nu_{z}$ is generated, as a homogeneous Young measure over $\mathbb{D}$, by taking a suitable diagonal subsequence $(\psi_{j,\lambda_{j}})$ of $\psi_{j,\lambda}(w)=\lambda^{-1}\bigl{(}u_{j}(z_{0}+\lambda w)-u_{j}(z_{0})\bigr{)}$, where $j\to\infty$ and $\lambda_{j}\to 0$, see [5, Theorem 2.8]. In particular, $\psi_{j,\lambda_{j}}\colon\mathbb{D}\to\mathbb{C}$ is a sequence of homeomorphisms such that $\psi_{j,\lambda_{j}}\to A_{z}$ weakly in $\mathrm{W}^{1,2}(\mathbb{D})$ and $\sup_{j}\|K_{\psi_{j,\lambda_{j}}}\|_{\mathrm{L}^{q}(\mathbb{D})}<\infty$. Then we are precisely in position to apply Proposition 1.4, whereby we find a sequence $(f_{j})$ of principal homeomorphisms of class $\mathrm{W}^{1,1}_{\operatorname{loc}}(\mathbb{C})$ satisfying $f_{j}\to A_{z}$ locally uniformly in $\mathbb{C}$, $f_{j}\to A_{z}$ weakly in $\mathrm{W}^{1,2}_{\operatorname{loc}}(\mathbb{D})$, $A_{f_{j}}=\fint_{\mathbb{D}}\!\textup{D}f_{j}\,\mathrm{d}m(z)\to A_{z}$, $K_{f_{j}}=K_{\psi_{j,\lambda_{j}}}$ a.e. in $\mathbb{D}$ and $(\textup{D}f_{j}|_{\mathbb{D}})$ generates $\nu_{z}$. Now as $f_{j}$ are principal maps it follows that $(J_{f_{j}})$ is bounded in $\mathrm{L}^{1}_{\operatorname{loc}}(\mathbb{C})$, and therefore we have for $s\in(\frac{p}{2-p},q)$ that
$|\textup{D}f_{j}|^{p}=J_{f_{j}}^{\frac{p}{2}}K_{f_{j}}^{\frac{p}{2}}\leq J_{f_{j}}^{\frac{ps}{2s-p}}+K_{f_{j}}^{s}$
is equiintegrable on $\mathbb{D}$. In view of Proposition 3.5 the sequence $\bigl{(}{\bf{E}}(z,\textup{D}f_{j}(w))\bigr{)}$ is equiintegrable over $\mathbb{D}$ and since ${\bf E}(z,\cdot)$ is principal quasiconvex, it is in particular rank-one convex on $\mathbb{R}^{2\times 2}_{+}$, and so continuous there, hence
$\int_{\mathbb{R}^{2\times 2}}\!{\bf E}(z,A)\,\mathrm{d}\nu_{z}(A)=\lim_{j\to\infty}\fint_{\mathbb{D}}\!{\bf E}(z,\textup{D}f_{j}(w))\,\mathrm{d}m(w)\geq\lim_{j\to\infty}{\bf E}(z,A_{f_{j}})={\bf E}(z,A_{z}),$